# Accurate RNA 3D structure prediction using a language model-based deep learning approach

### v2

@@ Methods
Our foundation model provides meaningful representations that are inferred from standalone sequence information. These representations may improve performances in various downstream tasks, especially for those with insufficient annotated data. Inspired by recent studies 48, 49, we leverage a general transformer architecture. In particular, our framework was built on the bidirectional transformer language model proposed in BERT devlin2018bert, followed by the unsupervised training scheme. We named our framework ‘RNA-FM’ as it represents a foundational model for future RNA-related studies (Supplementary Figure 2). Below, we detail how we constructed the large-scale non-coding RNA (ncRNA) dataset, followed by model and training details.
@@
Our RNA-FM framework comprises 12 transformer-encoder blocks, inspired by BERT devlin2018bert, 48. Each block includes a 640-hidden-size feed-forward layer and a multi-head self-attention layer with 20 heads, along with layer normalization and residual connections applied pre- and post-block, respectively. For an RNA sequence of length $L$, RNA-FM takes raw sequential tokens as input, mapping each nucleotide into a 640-dimensional vector via an embedding layer, forming an $L\times 640$ embedding matrix. This matrix passes through each encoder block, retaining its size throughout, and is followed by a softmax layer to predict corresponding tokens, including 16 nucleotides and four specific functional identifiers. Additional model details are in the Supplementary Information.
During pre-training, we employed self-supervised training akin to BERT devlin2018bert, randomly replacing 15% of nucleotide tokens with a special mask token. If the $i$-th token was chosen, it was replaced with (1) the [MASK] token 80% of the time, (2) a random token 10% of the time, and (3) left unchanged 10% of the time. We trained the model using masked language modeling (MLM) devlin2018bert, predicting the original masked token via cross-entropy loss. This training strategy is formulated as an objective function as follows:
@@
The secondary structure constraint loss, $L_{ss3d}$, encodes secondary structural information directly into 3D prediction. In order to unify the calculation of different types of base pairing constraints in 3D space, we introduced four fixed pseudo-atoms (T1, T2, T3, T4) in the local coordinate system of a base 50 (Supplementary Figure 1). $L_{ss3d}$ aims to constrain the pseudo atoms in two base-paired nucleobases to satisfy the base-pairing property (base-base interactions). For two residue $m$ and $n$, we computed the pairwise distance of the fixed points: $\mathbb{D}^{m,n}=\{d_{i,j}^{m,n}|i,j\in\{1,2,3,4\}\}$, where $m$, $n$ denote two RNA residues and $i$, $j$ are the indexes of the four atoms. We defined $L_{ss3d}$ as follows:
@@
As a preventive measure to resolve any remaining structural clashes and violations, we may relax our model predictions using a restrained energy minimization procedure, such as AMBER 51 and BRiQ 50. Specifically, we minimized the AMBER force field using harmonic restraints, allowing the system to maintain a close relationship with its input structure. This post-prediction relaxation also enforces the geometric features of phosphodiester bonds. Our empirical evidence indicates, as measured by RMSD and TM-score, that while this final relaxation does not improve the model’s accuracy, it eliminates distracting stereochemical violations without compromising accuracy.
@@ Acknowledgments
This work was supported by the Chinese University of Hong Kong (CUHK; award numbers 4937025, 4937026, 5501517, 5501329, 8601603, 8601663) and the Research Grants Council of the Hong Kong Special Administrative Region, China (Hong Kong SAR; project numbers CUHK 24204023, CUHK 14222922, and RGC GRF 2151185). Additional support was provided by the Innovation and Technology Commission of the Hong Kong SAR (project no. GHP/065/21SZ to Y. Li). J.W. was supported by a Hong Kong PhD Fellowship (award no. PF22-73180) from the Research Grants Council of the Hong Kong SAR, China, and an IdeaBooster Fund (project no. IDBF24ENG06) from CUHK. The work is part of the Antibiotics-AI Project, directed by J.J.C., and supported by the Audacious Project, Flu Lab, LLC, the Sea Grape Foundation, R. Zander, H. Wyss for the Wyss Foundation, and an anonymous donor. F.W. was supported by the National Institute of Allergy and Infectious Diseases of the NIH (award no. K25AI168451). D.L. acknowledges the startup funding from ASU. We thank Xiang-Jun Lu for guidance on inter-helical twist experiments.

### v3

@@ Methods
Our foundation model provides meaningful representations that are inferred from standalone sequence information. These representations may improve performances in various downstream tasks, especially for those with insufficient annotated data. Inspired by recent studies 48, 49, we leverage a general transformer architecture. In particular, our framework was built on the bidirectional transformer language model proposed in BERT 50, followed by the unsupervised training scheme. We named our framework ‘RNA-FM’ as it represents a foundational model for future RNA-related studies (Supplementary Figure 2). Below, we detail how we constructed the large-scale non-coding RNA (ncRNA) dataset, followed by model and training details.
@@
Our RNA-FM framework comprises 12 transformer-encoder blocks, inspired by BERT 50, 48. Each block includes a 640-hidden-size feed-forward layer and a multi-head self-attention layer with 20 heads, along with layer normalization and residual connections applied pre- and post-block, respectively. For an RNA sequence of length $L$, RNA-FM takes raw sequential tokens as input, mapping each nucleotide into a 640-dimensional vector via an embedding layer, forming an $L\times 640$ embedding matrix. This matrix passes through each encoder block, retaining its size throughout, and is followed by a softmax layer to predict corresponding tokens, including 16 nucleotides and four specific functional identifiers. Additional model details are in the Supplementary Information.
During pre-training, we employed self-supervised training akin to BERT 50, randomly replacing 15% of nucleotide tokens with a special mask token. If the $i$-th token was chosen, it was replaced with (1) the [MASK] token 80% of the time, (2) a random token 10% of the time, and (3) left unchanged 10% of the time. We trained the model using masked language modeling (MLM) 50, predicting the original masked token via cross-entropy loss. This training strategy is formulated as an objective function as follows:
@@
The secondary structure constraint loss, $L_{ss3d}$, encodes secondary structural information directly into 3D prediction. In order to unify the calculation of different types of base pairing constraints in 3D space, we introduced four fixed pseudo-atoms (T1, T2, T3, T4) in the local coordinate system of a base 51 (Supplementary Figure 1). $L_{ss3d}$ aims to constrain the pseudo atoms in two base-paired nucleobases to satisfy the base-pairing property (base-base interactions). For two residue $m$ and $n$, we computed the pairwise distance of the fixed points: $\mathbb{D}^{m,n}=\{d_{i,j}^{m,n}|i,j\in\{1,2,3,4\}\}$, where $m$, $n$ denote two RNA residues and $i$, $j$ are the indexes of the four atoms. We defined $L_{ss3d}$ as follows:
@@
As a preventive measure to resolve any remaining structural clashes and violations, we may relax our model predictions using a restrained energy minimization procedure, such as AMBER 52 and BRiQ 51. Specifically, we minimized the AMBER force field using harmonic restraints, allowing the system to maintain a close relationship with its input structure. This post-prediction relaxation also enforces the geometric features of phosphodiester bonds. Our empirical evidence indicates, as measured by RMSD and TM-score, that while this final relaxation does not improve the model’s accuracy, it eliminates distracting stereochemical violations without compromising accuracy.
@@ Acknowledgments
Funding: This work was supported by the Chinese University of Hong Kong (CUHK; award numbers 4937025, 4937026, 5501517, 5501329, 8601603, 8601663) and the Research Grants Council of the Hong Kong Special Administrative Region, China (Hong Kong SAR; project numbers CUHK 24204023, CUHK 14222922, and RGC GRF 2151185). Additional support was provided by the Innovation and Technology Commission of the Hong Kong SAR (project no. GHP/065/21SZ to Y. Li). J.W. was supported by a Hong Kong PhD Fellowship (award no. PF22-73180) from the Research Grants Council of the Hong Kong SAR, China, and an IdeaBooster Fund (project no. IDBF24ENG06) from CUHK. The work is part of the Antibiotics-AI Project, directed by J.J.C., and supported by the Audacious Project, Flu Lab, LLC, the Sea Grape Foundation, R. Zander, H. Wyss for the Wyss Foundation, and an anonymous donor. F.W. was supported by the National Institute of Allergy and Infectious Diseases of the NIH (award no. K25AI168451). D.L. acknowledges the startup funding from ASU. We thank Xiang-Jun Lu for guidance on inter-helical twist experiments.
Publication: We thank all our collaborators and the editorial team at Nature Methods for facilitating this publication. This work appears in Nature Methods 21, 2287–2298 (2024). doi:10.1038/s41592-024-02487-0